Sol Summary – November 22th

Sol: 9
Summary Title: Trouble on Mars
Author’s name: Jas Purewal
Mission Status: Nominal
Sol Activity Summary:
A full day of EVA’s today. Morning walking search and rescue EVA. Afternoon EVA to test the robotic arm and collect biological samples. Crew Biologist is working on documentation of samples. Collected 43 unique individual samples.
Look Ahead Plan: XO, HSO and Biologist EVA at 10:30am to the Special Region, making a stop at Gateway to Lith for a Biological sample collecting mission. Afternoon will be Thanksgiving celebrations.
Anomalies in work: None
Weather: Rather nice. Sunny, clear. No rain.
Crew Physical Status: One crew member still feels a little dizzy, but improved from yesterday.
EVA: EVA 10 – XO and HSO walked to the Gateway to Candor and simulated a twisted ankle. EVA crew radioed into HABCOM for help. HABCOM sent two crew in two rovers (Opportunity and Curiosity) to rescue the injured crew member. All crew returned safely to the hab at the expected EVA end time. A very successful rescue mission was completed. The second EVA of the day, EVA 11 in the afternoon, was to Cowboy Corner. The Commander, Biologist and Engineer took Spirit and Perseverance. 7 samples collected including ice, root soil and plants. Pleiades was 50% successfully collecting rocks, it broke as expected from the handle and the pads from the long arms. These were the weak points that were deliberately pushed to breaking point. As it was breaking, it was visually confirmed that the operation was being affected by the weather (e.g. accumulation of debris).
Reports to be filed:
Commander’s report
EVA report
EVA request
Operations report
Sol Summary report
Journalist report

Support Requested: none

Operations Report – November 22th

SOL: 9
Name of person filing report: Jas Purewal
Non-nominal systems: Had to fix the drain under the kitchen sink again.
Notes on non-nominal systems: NA
Spirit rover used: Yes
Hours: (before EVA) 199.9
Beginning charge: (Before EVA) 100%
Ending charge: (On return from EVA, before recharging) 76%
Currently charging: No
Opportunity rover used: Yes
Hours: 105.1
Beginning charge: 100%
Ending charge: 88%
Currently charging: Yes
Curiosity rover used: Yes
Hours: 212.5
Beginning charge: 100%
Ending charge: 91%
Currently charging: Yes
Perseverance rover used: Yes
Hours: 249.7
Beginning charge: 100%
Ending charge: 58%
Currently charging: No
General notes on rovers: Percy’s left rear tail light is out
Summary of Hab operations: put notes here
WATER USE: 25 gallons
Water (static tank): 175 gallons
Static tank pipe heater (on or off): on
Static tank heater (On or off) on
Toilet tank emptied: No
Summary of internet: Good
Summary of suits and radios: optimal
Summary of GreenHab operations: None
WATER USE: 6.5 gallons
Heater: On
Supplemental light: Off
Harvest: none
Summary of ScienceDome operations: None
Dual split: (Heat or AC, On or Off) Heat, Off
Summary of RAM operations: (Tools used, work done) Received first aid kit resupply
Summary of any observatory issues: NA
Summary of health and safety issues: One crew member still feels dizzy but improving.
Questions, concerns and requests to Mission Support: Whilst we are conserving water, at our current rate of use we predict that we will be down to our last 50 gallons by Sol 11.

Journalist Report – November 22th

Sol: 9
Title: Trouble on Mars

Author: Izabela Shopova, Crew Journalist

Sol 9 begins with an early EVA. Just before 9AM we are all downstairs helping the SHO and Crew Scientist to suit up. They are going on a walking exploratory EVA, but we have to troubleshoot some radio problems first. Jas appears to have a personal magnetic field – all radio sets stop working once she puts them on. By the time everything is tested and we are confident that comms work well, it’s almost 9:30AM. The day is clear and the EVA is going well until a distress call comes – Elizabeth reports that Jas has twisted her ankle on the steep slope and can’t walk back. Emergency response is initiated immediately – our crew engineer and crew biologist suit up, take two of the rovers, a first aid kit, a warm blanket, and some improvised materials to serve as a splinter if necessary. Commander and crew journalist stayed back in Hab for comms support. Luckily the injury occurred not too far from Hab where the EVA crew was within comms range. Everyone managed to get back in the airlock on time, just a couple of minutes later than the planned end of the EVA. When the patient is finally safe in the Hab, the truth is revealed – it was only a rescue exercise. We all exhale deeply with relief and cheer up. Jas walks around triumphantly and the crew gets into a lively debriefing session over lunch. Phew! This one went well.
Afternoon EVA was not as dramatic. The 3 crewmates tested successfully the mechanical arm Pleiades Atlas, discovered a desert oasis, a silently screaming rock, took amazing photos, and returned back to the Hab a few minutes earlier than planned The rest of the afternoon was quiet and productive – we all sat at our laptops and did as much work as possible while the sunset was setting the Martian rocks on fire and our crew engineer fixed the leaking sink again.

After the early dinner, we had a presentation on the benefits and the process of making yogurt. The main planning topic tonight was Thanksgiving – what will we cook, how will we celebrate, and how to fit the laborious preparations into our busy schedules.

Life on Mars is never boring.

GreenHab Report – November 22th

Sol 9
GreenHab Officer: Caitlyn Hubric
Environmental control: (heater, fan, door open?) heater on, door closed, 20% humidity
Average temperatures: 43.3ºC/13.5ºC
Hours of supplemental light: N/A
Daily water usage for crops: 24.6 L ≈ 6.50 gal
Daily water usage for research and/or other purposes: 0 gal
Water in Blue Tank ~251 gallons
Time(s) of watering for crops: twice, morning and night, 9am-5pm
Changes to crops:
-I replanted the assorted microgreens in the gray bin to the right of the shelves. I took the already open seed packet named “microgreens flavorful mix” and sprinkled the remaining seeds all over the planter.
-I noticed today that the gray bin to the right of the shelves was staying super wet and wasn’t draining properly after watering. After inspecting further, I realized the drainage hole was still screwed shut. I put one of the white hanging bins underneath it and opened the drainage hole. It is now draining properly and this should help prevent root rot in future plants.
Harvest: (include which crop and mass in grams) N/A
Support/supplies needed: N/A

EVA Report – November 22th

EVA # 10
Author: Elizabeth Balga
Purpose of EVA: Emergency response scenario – minor ankle injury.
Start time: 9:00 AM
End time: 11:00 AM
Narrative: XO and HSO initiated a walking EVA to the Gateway to Candor region by way of Cow Dung Rd. Upon arriving at the destination, the EVA team simulated a one-way communications failure with the habitat as well as an ankle injury. As the HABCOM for the EVA, the commander sent a rescue team consisting of the Crew Engineer and Crew Biologist to the Intersection of Cow Dung Rd and Gateway to Candor via rovers (Oppy & Curi). Two crewmembers — CDR and Communications Officer — remained at the habitat. The rescue team also brought a first aid kit, splinting materials, and a jacket for additional warmth. The HSO and rescue crew attempted to immobilize the ankle injury. All four crew members returned to the habitat via rover.
Destination: Gateway to Candor
Coordinates (use UTM WGS 84): 519000, 4252000
Participants: Initial team — XO, HSO; Rescue team — BIO, ENGR
Road(s) and routes per MDRS Map:
Along Cow Dung Road by foot
Mode of travel: Walking (Rescue team via rover)

EVA # 11
Author: Elizabeth Balga for Judith Marcos & Caitlyn Hubric
Purpose of EVA:
Primary: In-situ demonstration of anchor device carrying capacity
Secondary: Biological sample collection.
Start time: 2:00 PM
End time: 4:00 PM
Narrative: CDR, ENGR, and BIO completed an EVA to Cowboy Corner. ENGR prepared for and performed an engineering demonstration of the Pleiades instrument. During sample collection with the instrument, it broke in two locations — along the handle and along the arms — which informed the weak points of the device. The Crew Engineer was able to visually confirm that the device’s operation was affected by accumulation of debris, as suspected. BIO collected seven unique samples including ice, plants, and soil from the roots of said plants. Additionally, BIO noted upon return that Percy’s left rear tail light was out.
Destination: Cowboy Corner
Coordinates (use UTM WGS 84): 519000, 4253000
Participants: CDR, ENGR, BIO
Road(s) and routes per MDRS Map:
Along Cow Dung Road to Cowboy Corner by rover
Mode of travel: Rover, Walking (Percy, Spirit)

Commander Report – November 22th

Sol: 9
Summary Title: Trouble on Mars
Author’s name: Dr. Jenni Hesterman, Commander
Mission Status: Nominal
Commander Report:

The crew was up early today to discuss the day ahead. Sadly, we only have 3 more days on Mars and endeavor to maximize opportunities to hone skills for operating in a remote, austere environment with limited resources.

First up was a local walking EVA for the XO and Health and Safety Officer (HSO) to the Gateway to Candor area. Upon arrival, the crew notified the station that the XO injured her ankle and then lost communication. Two of the remaining four crew members in the station quickly donned their gear and entered the airlock within minutes to respond. They took two rovers and headed to the last known position of the crew with a first aid kit and a leg splint.

Once the crews made contact, they splinted the XO’s leg and she was loaded into a rover. After all crewmembers were safe in the station, the XO and HSO informed us this was a drill and we spent an hour discussing the strengths and weaknesses in our response.

After lunch, three crew members set out with two rovers on a biological mission to Cowboy Corner. We were surprised to find several areas with ice, despite the mild temperature and sunny day. The Crew Biologist took samples and photographed several species of lichens. We spotted an unreachable lagoon in the distance, surrounded by green plants and a tree. The Crew Engineer successfully accomplished a carry capacity test with the Pleiades Anchor, which will help astronauts easily retrieve rock and soil samples without having to bend over or squat. She deliberately pushed the robot to the breaking point to test its ability to function with an accumulation of debris and on different surfaces.

Following a dinner of cheesy pasta with a ham or tuna topping, the crew accomplished mission planning for the rest of the week. We have one long EVA tomorrow, then the American crewmembers will host an early Thanksgiving celebration for their international colleagues.

Copyright © The Mars Society. All rights reserved. | Main Site